I have two problems that I would love some help on. I am so confused.
Solve for x: 4x^2 = - 25x - 25

Solve by completing the square: m^2 - 49/2m = 1/2
Thank you in advance!!

Answer by edjones(8007)

4x^2 = - 25x - 25
4x^2+25x+25=0
4*25=100. Multiply a * c
5 and 20 are factors of 100 whose sum is 25.
4x^2+5x+20x+25=0
x(4x+5)+5(4x+5) factoring by grouping.
(x+5)(4x+5)=0
x=-5
4x+5=0, 4x=-5, x=-5/4
.
Ed